OpenStudy (anonymous):

Four friends compared there bike riding times. Jess rode 4.25 km in 20 min. Corey rode 3.75 km in 18 min. Mia rode 3.5 km in 17 min. Zoey rode 3.6 km in 16 min. Who rode at the greatest rate of speed? A. Corey B. Jess C. Mia D. Zoey

OpenStudy (anonymous):

Just divide the distances by the minutes of each person. Then see who has the lowest number.

OpenStudy (anonymous):

is it d im not sure

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

man!

OpenStudy (anonymous):

Did you divide the distances by their speeds?

OpenStudy (anonymous):

yes guess my calculator is dumb now or it me

OpenStudy (anonymous):

Which one has the lowest value?

OpenStudy (anonymous):

then its ...... Jess for me it came out d

OpenStudy (anonymous):

@texaschic101 Thanks for the medal, I guess.

OpenStudy (anonymous):

Not Jess.

OpenStudy (texaschic101):

it is asking for greatest rate of speed.....you sure it would be the lowest value ?

OpenStudy (anonymous):

Ohhhhh.... So It's asking for the slowest person?

OpenStudy (texaschic101):

its asking for the fastest person

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

greatest rate of speed

OpenStudy (texaschic101):

its asking who was the fastest

OpenStudy (anonymous):

ya thats what i thought

OpenStudy (anonymous):

Maybe now is the time to give direct answers maybe.... You think I should? @texaschic101

OpenStudy (anonymous):

or or u can tell me the right way how to do this !

OpenStudy (anonymous):

maybe

OpenStudy (texaschic101):

thats up to you....Moderators do not like me that much....I have been suspended for giving answers....just be careful.

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

that never happened with me

OpenStudy (anonymous):

The answer is..... is..... C. Am I right? @texaschic101

OpenStudy (texaschic101):

jesse -- 4.25/20 = .2125 corey -- 3.75/18 = .2083 mia -- 3.50/17 = .2058 zoey -- 3.6/16 = .225 fastest is zoey

OpenStudy (anonymous):

What?? But Mia has .2058. That's less. Or does it have to be the higher number?

OpenStudy (texaschic101):

it has to be the highest number...highest number is the fastest speed

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

I guess you were right in the beginning @lululife

OpenStudy (texaschic101):

I think it is zoey.....she traveled .225 km per min

OpenStudy (texaschic101):

and the greatest rate of speed means the fastest

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

so who should i give the medal to

OpenStudy (texaschic101):

my brain hurts now...thinking too much...lol

OpenStudy (texaschic101):

give it to AnswerMyQuestions

OpenStudy (anonymous):

ill give the medal to u then

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

no...give it to AnswerMyQuestions....he/she told you how to do it

OpenStudy (anonymous):

ya ok i will

OpenStudy (texaschic101):

cool :) all good

OpenStudy (anonymous):

Well, you were the one who corrected me. @lululife Would have gotten it wrong if you weren't here.

OpenStudy (texaschic101):

I am not worried about medals....I can't get any higher anyways

OpenStudy (anonymous):

There, gave you one.

OpenStudy (texaschic101):

thank you :)

OpenStudy (anonymous):

Np :D